b cross c whole square equals.

Grade:12

4 Answers

Aravind Bommera
11 years ago

b2+c2+2bc

VEERU YERRAYYA
11 years ago

|b×c|2 =|b|2|c|2 (sin@)2

lokesh soni
11 years ago

lbl2lcl2sin2θ

kabir pandey
6 years ago
 \left | aXb |^{^{2}}    i.e. vector a cross vector b, whole square equals to \left |a|^{^{2\left | b \right |^{2}sin^{2}\Theta.   it is the outcome of dot product of a cross b with a cross b.